Question 4 A 4.31 nC charge is placed at the origin ( 0.0,y0.0). What is the y-component of the electric field at the location (z -6.0 m, y 8.0 m) due to the charge at the origin?
Answer #1

4.)

Since, Electric field due to point charge(\underset{E}{\rightarrow}) = [k*q/(r)^3]*\underset{r}{\rightarrow}

where, k = 9*10^9

q = charge = 4.31 nC = 4.31*10^-9 C

\underset{r}{\rightarrow} = displacement vector = (6.0 - 0.0)i + (8.0 - 0.0)j

\underset{r}{\rightarrow} = 6.0 i + 8.0 j

So, distance(r) = |\underset{r}{\rightarrow}| = sqrt(6.0^2 + 8.0^2)

|\underset{r}{\rightarrow}| = 10.0 m

So,  \underset{E}{\rightarrow} = [(9*10^9)*(4.31*10^-9)/(10.0)^3]*(6.0 i + 8.0 j)

\underset{E}{\rightarrow} = 0.03879*(6.0) i + 0.03879*(8.0) j

\underset{E}{\rightarrow} = 0.233 i + 0.31 j

therefore y-component of electric field(Ey) = 0.31 N/C

Please upvote.